Let R1 and R2 be two relation on real number such that R1 = {(x, y)|x < y} and R2 =

{(x, y)|x > y}, then find R1 ∪ R2,R1 ∩ R2,R1 − R2, R2 − R1, and R1+LR2.

1
Expert's answer
2022-03-16T05:08:06-0400

R1"\cup" R2={(x,y)|x>y or x<y}={(x,y)|x!=y}

R1"\cap" R2={(x,y)|x>y and x<y}="\varnothing"

R1-R2=R1-R1"\cap" R2=R1 -{(x,y)|x>y and x<y}=R1-"\varnothing"=R1

R2-R1=R2-R1"\cap" R2=R2 -{(x,y)|x>y and x<y}=R2-"\varnothing"=R2

R1+LR2=R1"\cup" LR2-R1"\cap" LR2={(х,у)"|\forall x \in\R,\forall y \in \R"}-∅={(x,y)"|\forall x \in\R,\forall y \in \R" }